OpenStudy (anonymous):

Find dy/dx for y=Sqrt(x(x+7)) Using logarithmic differentiation. Answer can only be in terms of x. Thanks in advance for all the help yall have been giving me tonight! Learning much more than I do by going to class, lol!

OpenStudy (anonymous):

Where I get particularly confused is how do I eliminate the 1/y variable from the left side of the equation that comes from taking the derivative of (lny)' if that helps anyone understand my problem!

OpenStudy (anonymous):

No wait you can it's just tedious

OpenStudy (anonymous):

Just plug in (x^2 + 7x)^(1/2) for y after you derive

OpenStudy (anonymous):

And then put that back on the right side of the equation

OpenStudy (freckles):

\[y=\sqrt{x(x+7)} \\ \ln(y)=\ln([x(x+7)]^\frac{1}{2}) \\ \ln(y)=\frac{1}{2}(\ln(x(x+7)))\] It would be totally cute to just write that ln(x(x+7)) as ln(x)+ln(x+7)

OpenStudy (freckles):

can you tell us how far you have gotten

OpenStudy (freckles):

like have you gotten to the differentiating part if so can we see it

OpenStudy (anonymous):

I have y'/y=.5(2x+7/(x^2+7x)) I'm pretty sure that's correct? Did it a little different than you though I distributed the x(x+7) before I did anything

OpenStudy (freckles):

looks fabulous

OpenStudy (freckles):

now do what @master50777 said

OpenStudy (freckles):

multiply the y on both sides

OpenStudy (freckles):

and then replace y with sqrt(x^2+7x)

OpenStudy (anonymous):

OHH!!!!!!! It wasn't clicking with me where you guys were getting what y=... I'm silly hahaha. Thanks again @master50777 , and also to you @freckles !

OpenStudy (freckles):

yah a lot of people forget the initial equation and what it was called not just you

OpenStudy (anonymous):

Awesome to learn to look out for that now though!

OpenStudy (freckles):

you will see it some more in calculus if you find y'' given an implicit equation for y

OpenStudy (freckles):

\[\text{ example } x^2+y^2=4 \\ \text{ find } y'' \\ 2x+2yy'=0 \\ y'=\frac{-2x}{2y}=\frac{-x}{y} \\ y''=\frac{(-1)y-y'(-x)}{y^2} \\=\frac{-y+y'x}{y^2} \text{ first replace } y' \text{ with } \frac{-x}{y} \\ y''=\frac{-y-\frac{-x}{y}(-x)}{y^2} =\frac{-y-\frac{x^2}{y}}{y^2} \text{ multiply both \top and bottom by } y \\ = \frac{-y^2-x^2}{y^3} =-\frac{x^2+y^2}{y^3} \\ \text{ but see initial equation } x^2+y^2=4 \text{ so we have } \\ y''=\frac{-4}{y^3}\] just for fun you can look at this or not

OpenStudy (anonymous):

No I'll definitely look over that! My teacher doesn't quite explain in her lectures the way I need and I thank you guys a ton this is saving my retrice the morn I learn the better!

OpenStudy (freckles):

i bet you will assume be covering that topic if you haven't already

OpenStudy (freckles):

soon not assume I'm sorry my first language is english but I don't like it sometimes :p

OpenStudy (anonymous):

more** not morn. Your English is perfect pal